Today you buy 1,000 shares of in Exxon Mobil Corp. at $66.59 per share. After one quarter, the share price rises to $83.78 per share. At the end of this quarter, Exxon Mobil pays a cash dividend of $1.45 per share. Assume that you reinvest your cash dividends in Exxon Mobil stock.

What is the percent increase in the number of shares that you hold? (Assume that fractional shares are available.)

Do not round values at intermediate steps in your calculations. Enter your answer in percent to two decimal places, but omit the % symbol. For example, enter 5.38% as 5.38 in your answer.
Answer #1

Total dividend received = 1.45*1,000 = $1,450

Price per share at the end of quarter = $83.78

Number of shares bought from dividend = 1,450/83.78

= 17.307233 shares

Percent increase in number of shares = 17.307233/1,000

= 1.73%

i.e. 1.73